# The Impact of Coreset Selection on Spurious Correlations and Group Robustness

### v1

@@ Abstract
Coreset selection methods have shown promise in reducing the training data size while maintaining model performance for data-efficient machine learning. However, as many datasets suffer from biases that cause models to learn spurious correlations instead of causal features, it is important to understand whether and how dataset reduction methods may perpetuate, amplify, or mitigate these biases. In this work, we conduct the first comprehensive analysis of the implications of data selection on the spurious bias levels of the selected coresets and the robustness of downstream models trained on them. We use an extensive experimental setting spanning ten different spurious correlations benchmarks, five score metrics to characterize sample importance/ difficulty, and five data selection policies across a broad range of coreset sizes. Thereby, we unravel a series of nontrivial nuances in interactions between sample difficulty and bias alignment, as well as dataset bias and resultant model robustness. For example, we find that selecting coresets using embedding-based sample characterization scores runs a comparatively lower risk of inadvertently exacerbating bias than selecting using characterizations based on learning dynamics. Most importantly, our analysis reveals that although some coreset selection methods could achieve lower bias levels by prioritizing difficult samples, they do not reliably guarantee downstream robustness.
 [^1] Our results and code are available here
@@ 1 Introduction
We begin with a quick preview of some of our results to demonstrate these potential concerns. Here we consider the Waterbirds [20] dataset, where labels (waterbirds vs landbirds) co-occur with the backgrounds (land, water), forming four distinct groups. In the training set, $95\%$ of the waterbirds appear on water backgrounds, and $95\%$ of landbirds appear on land backgrounds, making the backgrounds a spurious feature. This spurious association can lead the model to predict the backgrounds rather than the birds, incurring disproportionately higher errors for waterbirds on land, and landbirds on water [17, 28]. In Figure 1 Left, we evaluate a classifier trained on small random subsets of the dataset. Uniform random sampling (or random coreset selection) is considered a strong coreset baseline [11, 10]. We consider both average accuracy on the full dataset as well as worst-group accuracy (the lowest accuracy among the four subgroups [20]) to evaluate the group-robustness of the classifier. Although classifiers continue to achieve high average accuracy with very small coresets, their robustness as measured by worst-group accuracy declines substantially. Most coreset selection methods first assign a score for each sample to characterize a notion of difficulty. They subsequently select a coreset by prioritizing a part of this score distribution as defined by a selection policy. In Figure 1 Right we plot the distribution of average and worst-group accuracies for models trained on 10% coresets selected using different combinations of sample characterization scores [27, 15, 14, 13, 12] and selection policies [11, 12, 13]. We show that different coresets with comparable downstream average accuracies can result in drastically different levels of group-robustness. Hence, we emphasize the crucial need for a deeper evaluation to understand how coreset selection impacts model robustness and how to select coresets in the presence of spurious correlations.
@@ 2 Related Work
Recent work uncovered that the optimal policy for selecting coresets is deeply nuanced by the dataset size and the coreset size: selecting the hardest samples can be suboptimal for learning an accurate classifier, especially when the coreset size is very small [13]. Furthermore, there seem to be many sources of hardness/difficulty for samples [40], and not all of them convey important information to train generalizable classifiers. Given that many sample characterization scores are strongly correlated [41], we investigate whether the concept of difficulty used in data selection is tied to spurious correlations.
@@
Data selection meets group-robustness. Pruning out instances of certain groups or classes when datasets are imbalanced, also coined "subsampling," aims to reduce bias and improve robustness of the downstream classifiers [54]. This is in line with previous work on data interventions for spurious correlation mitigation. A recent study [55] hints that data pruning may mitigate distributional bias in trained models- however it was limited to one data selection method: EL2N [27]. Distributionally robust data pruning (DRoP) [56] emulates this behavior at the class-level, retaining more samples from the more difficult classes. However, we reiterate the difficulty of our setting compared to this, since group labels are often unavailable, and there could be unknown spurious correlations. Therefore, the specific implications of data selection when datasets contain hidden biases remain underexplored.
@@ 3 Experimental setup
We consider several sample selection policies including: the Diff (Difficult) policy, which selects the highest scoring samples for the coreset, the Eas (Easy) policy, which selects the lowest scoring samples [13], the Med (Median) policy, which selects samples closest to the median of the score distribution [12], and the Strat (Stratified) policy [11], which construct 50 bins ranking from lowest to highest scores and sample randomly from each bin based on the data budget allocated for each bin. We implement the R (Random) selection policy as a baseline. Additionally, we implement an oracle selection policy R-Gbal (Random-Groupbalanced) which picks an equal number of random samples from each group to create a group-balanced coreset. (The extent of group-balancing is limited by the number of total samples present from each group, since we do not oversample.) R-Gbal acts as an upper bound since it utilizes both class labels and spurious feature labels to identify the groups, whereas we assume that only class labels are available in our setting.
@@
We keep the number of training iterations fixed across our experiments by setting the number of training epochs to $N/s$ where $N$ is the number of training epochs on the full dataset, and $s$ is the coreset size as a fraction of the full dataset. We use models initialized with Imagenet-1K [26] pretrained weights for the visual datasets: ResNet18 [62] for cMNIST, and ResNet50 for others. NLP datasets were classified using BERT [63] model pretrained on Book Corpus and English Wikipedia data. The same models and initializations were used as surrogate models to calculate characterization scores. (Please refer Appendix B for complete training details.)
We evaluate the average performance of the downstream classifier using Average accuracy, which assumes the test set has a similar group distribution as the original training set. This is done by computing group accuracies and re-weighting them according to their proportions in the original train dataset [20, 46]. Lastly and most importantly, we measure the group-robustness of the downstream classifier using worst-group accuracy, which is the minimum of the individual group accuracies. A high worst-group accuracy would reflect that the model has learnt a solution that is robust to the known spurious correlation of the dataset.
@@ 4 Empirical analysis
4.1 Embedding-based characterization scores run a lower risk of inadvertently exacerbating bias compared to learning-based characterizations
@@
In more detail, two learning-based scoring methods, EL2N [27] and Uncertainty [14] are substantially better than the random baseline at classifying bias-conflicting vs bias-aligning samples across all 5 datasets; suggesting that coresets selected using these scores run the risk of systematically excluding certain groups, inadvertently exacerbating bias levels. The third learning-based scoring method, Forgetting [15], is more nuanced, appearing to contain bias information on Waterbirds (23.9% AP vs 4.7% AP for the random baseline) and Urbancars-C (13.8% vs 5.0%) but comparatively little on the other datasets (e.g., on MetaShift 35.9% AP vs 30.4% AP for the random baseline).
@@
Learning-based scores, on the other hand, unsurprisingly provide stronger indicators for bias, and in return, the bias levels of the resultant coresets vary largely with the executed selection policy. This can be seen in the left column of Figure 3 (which we will discuss in more detail in the next section); for example, selecting 10% of the Urbancars-C data with learning-based EL2N scores varies between bias level of 1.54 using Difficult, and 2.0 using Easy compared to the bias level of 1.92 on the random coreset. Therefore, care needs to be given when using learning-based coreset selection scores, especially in the presence of unknown spurious biases in the data. Hence, we focus more on the inadvertent risks of learning based selections in the subsequent sections.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under grant no 2112562 and 2145198 to O.R. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. All experiments, data collection, and processing activities were conducted at Princeton University. Meta was involved solely in an advisory role and no experiments, data collection or processing activities were conducted on Meta infrastructure.
@@ C.2 Coreset bias level is not a consistent indicator of downstream robustness
Figure 7: Data bias and classifier accuracies for Difficult (highest-scoring) and Easy (lowest-scoring) scoring samples using EL2N scores. Selecting the Difficult samples typically results in less biased coresets and corresponding more robust (highest worst-group accuracy) classifiers than Easy samples. The Difficult samples also tend to be more robust than those with Random selection. However, as coreset size gets smaller, we see a significant drop in average and worst-group accuracies for Difficult samples.

### v2

@@ Abstract
Coreset selection methods have shown promise in reducing the training data size while maintaining model performance for data-efficient machine learning. However, as many datasets suffer from biases that cause models to learn spurious correlations instead of causal features, it is important to understand whether and how dataset reduction methods may perpetuate, amplify, or mitigate these biases. In this work, we conduct the first comprehensive analysis of the implications of data selection on the spurious bias levels of the selected coresets and the robustness of downstream models trained on them. We use an extensive experimental setting spanning ten different spurious correlations benchmarks, five score metrics to characterize sample importance/ difficulty, and five data selection policies across a broad range of coreset sizes. Thereby, we unravel a series of nontrivial nuances in interactions between sample difficulty and bias alignment, as well as dataset bias and resultant model robustness. For example, we find that selecting coresets using embedding-based sample characterization scores runs a comparatively lower risk of inadvertently exacerbating bias than selecting using characterizations based on learning dynamics. Most importantly, our analysis reveals that although some coreset selection methods could achieve lower bias levels by prioritizing difficult samples, they do not reliably guarantee downstream robustness. [^1] See https://github.com/princetonvisualai/Robustness-impacts-of-coreset-selection for results and code.
@@ 1 Introduction
We begin with a quick preview of some of our results to demonstrate these potential concerns. Here we consider the Waterbirds [20] dataset, where labels (waterbirds vs landbirds) co-occur with the backgrounds (land or water), forming four distinct subgroups. In the training set, $95\%$ of the waterbirds appear on water backgrounds, and $95\%$ of landbirds appear on land backgrounds, making the backgrounds a spurious feature. This association can lead the model to predict based on the background rather than on the bird, incurring disproportionately high errors for waterbirds on land, and landbirds on water [17, 28].
In Figure 1 Left, we demonstrate this disproportionate effect as the size of the coreset changes. We compare the worst-group accuracy (the lowest accuracy among the four subgroups [20]) and the average test accuracy of a model trained on coresets of different selection ratios. The coresets are selected using uniform random sampling (or random coreset selection), which is considered a strong baseline [11, 10]. Although classifiers continue to achieve high average accuracy with very small coresets, their robustness as measured by worst-group accuracy declines substantially. Hence, we highlight the importance of understanding how different coreset sizes impact the group robustness of resulting models.
In Figure 1 Right we conside more complex coreset selection methods and plot the distribution of average and worst-group accuracies for models trained on 10% coresets. Each point corresponds to the combination of one sample characterization score [27, 15, 14, 13, 12] and a selection policy [11, 12, 13]. We show that different coresets with comparable downstream average accuracies can result in drastically different levels of group-robustness. Hence, we emphasize the crucial need for a deeper evaluation to understand how different strategies for coreset selection impact model robustness and how to select coresets in the presence of spurious correlations.
@@ 2 Related Work
Recent work uncovered that the optimal policy for selecting coresets depends on the dataset size and the coreset size: selecting the hardest samples can be suboptimal for learning an accurate classifier, especially when the coreset size is very small [13]. Furthermore, there seem to be many sources of hardness/difficulty for samples [40], and not all of them convey important information for training generalizable classifiers. Given that many sample characterization scores are strongly correlated [41], we investigate whether the concept of difficulty used in data selection is tied to spurious correlations.
@@
Data selection meets group-robustness. Pruning out instances of certain groups or classes when datasets are imbalanced, also coined "subsampling," aims to reduce bias and improve robustness of the downstream classifiers [54]. This is in line with previous work on data interventions for spurious correlation mitigation. A recent study [55] hints that data pruning may mitigate distributional bias in trained models- however it was limited to one data selection method: EL2N [27]. A concurrent work [56] has explored how to prune out certain training samples to mitigate spurious correlations, in a setting where the spurious signal is relatively weaker. Distributionally robust data pruning (DRoP) [57] emulates this behavior at the class-level, retaining more samples from the more difficult classes. However, we reiterate the difficulty of our setting compared to this, since group labels are often unavailable, and there could be unknown spurious correlations.
@@ 3 Experimental setup
We consider several sample selection policies including: the Difficult policy, which selects the highest scoring samples for the coreset, the Easy policy, which selects the lowest scoring samples [13], the Median policy, which selects samples closest to the median of the score distribution [12], and the Stratified policy [11], which construct 50 bins ranking from lowest to highest scores and sample randomly from each bin based on the data budget allocated for each bin. We implement the Random selection policy as a baseline. Additionally, we implement an oracle selection policy Random-Groupbalanced which picks an equal number of random samples from each group to create a group-balanced coreset. (The extent of group-balancing is limited by the number of total samples present from each group, since we do not oversample.) R-Gbal acts as an upper bound since it utilizes both class labels and spurious feature labels to identify the groups, whereas we assume that only class labels are available in our setting.
@@
We keep the number of training iterations fixed across our experiments by setting the number of training epochs to $N/s$ where $N$ is the number of training epochs on the full dataset, and $s$ is the coreset size as a fraction of the full dataset. We use models initialized with Imagenet-1K [26] pretrained weights for the visual datasets: ResNet18 [63] for cMNIST, and ResNet50 for others. NLP datasets were classified using BERT [64] model pretrained on Book Corpus and English Wikipedia data. The same models and initializations were used as surrogate models to calculate characterization scores. (Please refer to Appendix B for complete training details.)
We evaluate the average performance of the downstream classifier using Average accuracy, which assumes the test set has a similar group distribution as the original training set. This is done by computing group accuracies and re-weighting them according to their proportions in the original train dataset [20, 46]. Lastly and most importantly, we measure the group-robustness of the downstream classifier using worst-group accuracy, which is the minimum of the individual group accuracies. A high worst-group accuracy reflects that the model has learnt a solution that is robust to the known spurious correlation of the dataset.
@@ 4 Empirical analysis
4.1 Embedding-based characterization scores run a lower risk of inadvertently exacerbating bias compared to learning-based scores
@@
In more detail, two learning-based scoring methods, EL2N [27] and Uncertainty [14] are substantially better than the random baseline at classifying bias-conflicting vs bias-aligning samples across all 5 datasets, suggesting that coresets selected using these scores run the risk of systematically excluding certain groups, inadvertently exacerbating bias levels. The third learning-based scoring method, Forgetting [15], is more nuanced, appearing to contain bias information on Waterbirds (23.9% AP vs 4.7% AP for the random baseline) and Urbancars-C (13.8% vs 5.0%) but comparatively little on the other datasets (e.g., on MetaShift 35.9% AP vs 30.4% AP for the random baseline).
@@
Learning-based scores, on the other hand, unsurprisingly provide stronger indicators for bias, and in return, the bias levels of the resultant coresets vary largely with the executed selection policy. This can be seen in the left column of Figure 3 (which we will discuss in more detail in the next section); for example, selecting 10% of the Urbancars-C data with learning-based EL2N scores varies between bias level of 1.5 using Difficult, and 2.0 using Easy compared to the bias level of 1.9 on the random coreset. Therefore, care needs to be given when using learning-based coreset selection scores, especially in the presence of unknown spurious biases in the data. Hence, we focus more on the inadvertent risks of learning based selections in the subsequent sections.
@@ 5 Conclusions
Algorithm 1 Coreset Selection ensuring Group-Robustness (Group information is unavailable)
Given: A training + validation set with class labels but no spurious annotations, and a desired coreset size range $N_{\min},N_{\max}]$ (dictated by application constraints).
Goal: Find the smallest coreset likely to yield high accuracy and group robustness on a test set with the same target classes and unknown group labels.
Method:
Initialize: Set coreset size $N=N_{\min}$.
while $N\leq N_{\max}$ do
  Select two coresets: $S_{N,\text{difficult}}$ using the difficult policy, $S_{N,\text{random}}$ using the random policy
  Train models: $M_{N,\text{difficult}}$ on $S_{N,\text{difficult}}$, $M_{N,\text{random}}$ on $S_{N,\text{random}}$
  if $\text{avg accuracy}(M_{N,\text{difficult}})\approx\text{avg accuracy}(M_{N,\text{random}})$ then
   return $S_{N,\text{difficult}}$ $\triangleright$ NN is “sufficiently large,” difficult coreset likely yeilds high robustness.
  else
   Increase NN $\triangleright$ NN is in the “small-data regime.” $S_{N,\text{difficult}}$ may not ensure robustness.
  end if
end while
[…]
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grants No 2112562 and 2145198. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. All experiments, data collection, and processing activities were conducted at Princeton University. Meta was involved solely in an advisory role and no experiments, data collection or processing activities were conducted on Meta infrastructure. We thank Angelina Wang for insightful discussions during the early stages of the project.
@@
  • C.4 - Further exploration into difficult coresets in small data regime
@@ C.2 Coreset bias level is not a consistent indicator of downstream robustness
Extended results for all scoring methods and all selection policies for 40% selection rate is shown in Table 5
Figure 7: Data bias and classifier accuracies for different selection policies using EL2N scores. Selecting the Difficult samples typically results in less biased coresets and corresponding more robust (highest worst-group accuracy) classifiers than Easy samples at higher selection rates. The Difficult samples also lead to more robust models than Random selection. However, as coreset size gets smaller, we see a significant drop in average and worst-group accuracies for Difficult samples. In such settings, Stratified and Median policies, which are consistently more biased than Difficult, counter-intuitively yield higher robustness.
@@
Dataset Baselines EL2N [27] Uncertainty [14] Forgetting [15] SelfSup [13] SupProto [12]
R R-gbal Diff Strat Med Eas Diff Strat Med Eas Diff Strat Med Eas Diff Strat Med Eas Diff Strat Med Eas
C-Mn 62.0 84.4 87.72 74.0 0.0 0.0 37.8 82.3 1.0 0.0 9.9 4.0 0.0 0.0 83.0 83.5 44.0 0.0 87.0 48.5 0.0 0.0
WB 58.2 83.3 74.6 73.1 50.5 30.2 72.6 71.2 51.0 28.2 75.4 79.6 53.8 52.0 50.9 51.0 69.2 37.1 66.5 58.8 59.4 30.0
UC-c 52.0 61.6 54.0 51.2 46.4 49.6 54.8 55.2 46.4 50.4 51.2 54 31.6 31.6 42.0 48.4 54.4 44.0 46.8 50.4 50.0 44.8
MSh 73.8 78.5 78.5 63.1 66.5 58.6 73.8 69.2 70.2 65.4 71.7 76.4 46.6 48.7 73.3 69.2 74.3 56.9 67.0 73.8 72.3 58.5
$C$ 77.4 72.0 63.0 51.7 63.9 43.2 60.2 64.2 61.0 75.2 60.9 65.3 82.0 79.5 77.8 78.4 70.3 79.3 76.2 77.1 74.1 75.2
Nic-s 44.0 40.0 44.0 44.0 34.0 32.0 44.0 42.0 36.0 30.0 44.0 40.0 46.0 40.0 44.0 16.0 34.0 36.0 38.0 42.0 42.0 38.0
UC-b 44.8 64.0 52.0 44.8 26.4 15.6 52.8 53.2 26.0 15.6 47.6 50.8 54.0 54.0 43.2 48.0 41.6 26.4 53.2 52.4 36.8 13.6
UC-a 16.8 50.4 23.2 19.2 6.4 7.2 22.4 26.4 8 7.2 20.0 23.2 16.0 16.0 11.2 17.6 20.8 9.6 20.8 19.2 17.6 5.6
MNL 60.5 57.6 46.0 65.6 61.8 46.0 61.5 64.0 53.8 49.2 59.5 57.8 52.4 52.4 55.2 60.8 55.7 58.3 - 63.9 62.6 55.4
Cel-h 58.9 71.1 38.3 47.8 68.9 79.0 42.2 42.2 62.8 78.9 27.2 42.8 64.4 64.4 41.7 66.1 63.3 78.9 50.0 59.4 63.9 82.2
Table 5: Worst-group accuracies of different selection policies within each scoring methods (learning-based: EL2N [27], Uncertainty [14], Forgetting [15], and embedding-based: SelfSup [13], SupProto [12]) at 40% selection rate. The highest worst-group-accuracies within each scoring method for each dataset are bolded.
@@ C.3 Trading off most difficult bias-conflicting samples to improve robustness
C.4 Further exploration into catastrophic loss of accuracy for difficult coresets in small data regime
We inspect train-test accuracy gaps for models trained on small-data regime coresets and observe that all models achieve 100% training accuracy. The drop in average accuracy for difficult coresets in the small data regime therefore suggests a high level of overfitting compared to other policies, also consistent with claims of Sorscher et al [13] as we discussed in Section 4.2.
It has been hypothesized that the catastrophic drop in accuracy for difficult coresets could be due to the coreset achieving low coverage on the data space. Zhen et al. [11] utilized the concept of p-partial r-cover to quantify this phenomenon, where r is some radius around each data of the coreset and p is the proportion of training data covered. We used their p-partial r-cover as a metric to investigate whether this pattern persists in datasets with strong spurious correlations. Using the features of ImageNet pre-trained ResNet-50, we selected a radius r where it would cover 95% of the training data. Using the r, we obtained the measured p from the selected coresets shown in Table 6. Significant decrease in p in difficult selected coresets at 5% and 2% across the datasets confirms the drop in coverage.
Selection Random Difficult Easy
Rate (baseline)
2% 80.3 45.4 82.3
5% 85.7 58.8 86.1
20% 93.3 92.4 90.9
(a) Waterbirds [20]
Selection Random Difficult Easy
Rate (baseline)
2% 61.0 7.9 71.1
5% 74.7 25.0 77.4
20% 89.6 80.0 88.4
(b) Metashift [31]
Table 6: p-partial r-cover achieved by Difficult and Easy selection policies for Waterbirds [20] and Metashift [31] datasets. At low selection rates, Difficult selection yeilds to significantly less coverage than Easy of Random selections (bolded).